The length of a rectangle is 3 more than its width. The perimeter of the rectangle is 58 centimeters. What are the rectangle's dimensions?
Directions: Write a system of equations to represent the situation. Then solve the system by substitution.

Please help me!!

Answer by mananth(16946)

let width be x cm
length will be x+3 cm
..
Perimeter = 2*(L+W)
--------------
2*(x+x+3)=58
--------------
2*(2x+3)=58
/2
(2x+3)=29
-3
2x+3-3=29-3
2x=26
/2
x=13 ----- the width
length = x+3=13+3=16 cm.
..
m.ananth@hotmail.ca